OpenStudy (anonymous):

A 20-foot ladder is leaning against a vertical wall. If the bottom of the ladder is being pulled away from the wall at the rate of 10 feet per second, at what rate is the area of the triangle formed by the wall, the ground, and the ladder changing, in square feet per second, at the instant the bottom of the ladder is 12 feet from the wall?

OpenStudy (anonymous):

PLEASSSSSE HELP

OpenStudy (amistre64):

the rate of change of area eh

OpenStudy (amistre64):

Whats our formula for the area of a rt triangle?

OpenStudy (anonymous):

a = 1/2 b*h

OpenStudy (amistre64):

good, and when we derive it implicitly with respect to time? we get?

OpenStudy (anonymous):

im not sure....is it 1/2 b*h ^-1/2

OpenStudy (amistre64):

\[A = \frac{1}{2}bh\] \[A' = \frac{1}{2}b'h+\frac{1}{2}bh'\] its just the product rule for derivatives

OpenStudy (anonymous):

oh i see

OpenStudy (amistre64):

but we can clean it up if we relate h in terms of b since it doesnt give us a rate of change for h does it

OpenStudy (anonymous):

so wat is h in terms of b

OpenStudy (amistre64):

it might be a trig function, which we might wanna stay away from; OR we could devise a scheme to find h' from b' ...

OpenStudy (amistre64):

i got it, pythag

OpenStudy (anonymous):

or h=16

OpenStudy (amistre64):

\[20^2 = h^2 + b^2\] \[20^2-b^2 = h^2\] derived we get \[-2b\ b' = 2h\ h'\] \[\frac{-2b\ b'}{2h} = h'\] \[\frac{-b\ b'}{h} = h'\]

OpenStudy (amistre64):

this is doable

OpenStudy (amistre64):

\[A' = \frac{1}{2}b'h+\frac{1}{2}bh'\] \[A' = \frac{1}{2}b'h+\frac{1}{2}b(\frac{-b\ b'}{h})\] messy, but doable; now we are told b' = 10, and b = 12 the pythag itself will give us our h at this moment in time and then its just plug and play

OpenStudy (anonymous):

wat do i plug for h

OpenStudy (amistre64):

\[20^2 = 12^2 + h^2\] \[400-144 = h^2\] \[\sqrt{256} = h\] \[16 = h\] right?

OpenStudy (anonymous):

ys

OpenStudy (amistre64):

then all together we get: \[A' = \frac{1}{2}b'h+\frac{1}{2}b(\frac{-b\ b'}{h})\] \[A' = \frac{1}{2}10(16)+\frac{1}{2}12(\frac{-12(10)}{16})\] \[A' = 80+6(\frac{-12(10)}{16})\] \[A' = 80+(\frac{-2(10)}{16})\] \[A' = 80-\frac{10}{8}\]

OpenStudy (anonymous):

the aswer choices are a) -70, b)35/2 c) 70 d) -35 e) 35

OpenStudy (anonymous):

wen u solve for wat u did isnt teh asnwer 78.75

OpenStudy (amistre64):

lets look over our math to be sure

OpenStudy (amistre64):

i get 35 when I do the calculations right

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

wat didu do wrng in the begining

OpenStudy (amistre64):

\[A' = \frac{1}{2}10(16)+\frac{1}{2}12(\frac{-12(10)}{16})\] \[A' = 10(8)+12(\frac{-12(5)}{16})\] \[A' = 80+3(-3(5))\] \[A' = 80-45=35\]

OpenStudy (amistre64):

the formulas were good, I just forgot how to multiply and add lol